OpenStudy (anonymous):

Line AB contains points A (1, 2) and B (−2, 6) The slope of line AB is zero undefined positive negative

OpenStudy (afifarahman):

\[slope = \frac{ y _{2} -y _{1} }{x _{2}-x _{1} }\]

OpenStudy (anonymous):

so what is the answer

OpenStudy (afifarahman):

\[x _{1} = 1 , y _{1}= 2\]

OpenStudy (afifarahman):

the same goes for x2 and y2

OpenStudy (anonymous):

and do these for too The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD? Line CD contains points A (4, 6) and B (−2, 6). The slope of line CD is Line QR contains (2, 8) and (3, 10) Line ST contains points (0, 6) and (−2, 2). Lines QR and ST are The equation of line QR is y = negative 1 over 2x + 1. Write an equation of a line perpendicular to line QR in slope-intercept form that contains point (5, 6). The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD? 1 over 2 2 negative 1 over 2 −2

OpenStudy (mrnood):

@VultureKV The answer is What YOU come up wit if you apply the equation which @afifaRahman kindly gave you above Put the values for x1,x2,y1 & y2 into that equation nd theanswer will be yours!

OpenStudy (anonymous):

06.02 Slope Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Multiple Choice Worth 4 points) (06.02 MC) Line QR contains (2, 8) and (3, 10) Line ST contains points (0, 6) and (−2, 2). Lines QR and ST are parallel because the product of the slopes is −1 perpendicular because the product of the slopes is −1 parallel because the slopes are the same perpendicular because the slopes are the same Question 2 (Multiple Choice Worth 4 points) (06.02 LC) The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD? 1 over 2 2 negative 1 over 2 −2 Question 3 (Multiple Choice Worth 4 points) (06.02 LC) Line CD contains points A (4, 6) and B (−2, 6). The slope of line CD is zero undefined positive negative Question 4 (Multiple Choice Worth 4 points) (06.02 MC) The equation of line QR is y = negative 1 over 2x + 1. Write an equation of a line perpendicular to line QR in slope-intercept form that contains point (5, 6). y = 2x + 16 y = negative 1 over 2x + 17 over 2 y = − 1 over 2x + 7 over 2 y = 2x − 4 Question 5 (Multiple Choice Worth 4 points) (06.02 MC) The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). y = −2x + 13 y = negative 1 over 2x + 7 y = 1 over 2x + 3 y = − 2x − 3

OpenStudy (mrnood):

http://openstudy.com/code-of-conduct PLease do NOT ask for answers THe help you get here will aid oyu in getting the answer for YOURSELF

OpenStudy (anonymous):

i need help

OpenStudy (mrnood):

You have had help - now try and do the answer Post your working her and you may get more help The Answer is NOT help - you will know no more having been given the answer than you did before

OpenStudy (anonymous):

i got the answer to it i mean i need the other ones done to help me

OpenStudy (mrnood):

Start with your first question, and put the values into the equation given above...

OpenStudy (mrnood):

What was your answer - and how did you get to it?

OpenStudy (anonymous):

i just cant seem to figure out the last ones :(

OpenStudy (mrnood):

Show me that you have done some work on the FIRST one and I will help Show the calculation for the slope of the first line

OpenStudy (anonymous):

i got y = negative 1 over 2x + 17 over 2 for number 4

OpenStudy (mrnood):

Show me your working for the first one